# Augmenting NER Datasets with LLMs: Towards Automated and Refined Annotation

### v1

@@ Appendix B Datasets
CoNLL03 (Tjong Kim Sang & De Meulder, 2003):
The CoNLL-2003 dataset is designed to recognize four types of named entities, including person names, place names, organization names, and other miscellaneous entities. The data file consists of four columns: word, part-of-speech tag, syntactic chunk tag, and named entity tag, with each item separated by a space. The named entity and syntactic chunk tags use the I-TYPE format  [^3] This tag means ”Inside” and is assigned to words that are within an entity of a specific type (TYPE). That is, this tag is used when a word is part of a specific entity and indicates the type of that entity. For example, ”I-PER” indicates that the word is part of a person’s name (Person)., and the B-TYPE  [^4] The B in B-TYPE means ”Beginning” and is the tag given to the first word of the second entity when directly followed by an entity of the same type. This tag is used to indicate that a new entity is beginning and to distinguish it from the previous entity. For example, ”B-PER” means that the new person’s name begins with this word. tag is introduced to indicate the start of a new phrase.
WikiGold (Balasuriya et al., 2009):
The Wikigold dataset is a manually annotated corpus for named entity recognition consisting of a small sample of Wikipedia articles. Words are labeled for each of the four named entity classes (LOC, MISC, ORG, and PER) of CONLL-03 described above.

### v2

@@ A.2 BERT Configurations:
A.3 Code Implementation
The experimental code for this study was implemented using Copilot. [^3] https://github.com/features/copilot
@@ Appendix B Datasets
B.1 CoNLL03:
The CoNLL-2003 dataset (Tjong Kim Sang and De Meulder, 2003) is designed to recognize four types of named entities, including person names, place names, organization names, and other miscellaneous entities. The data file consists of four columns: word, part-of-speech tag, syntactic chunk tag, and named entity tag, with each item separated by a space. The named entity and syntactic chunk tags use the I-TYPE format  [^4] This tag means ”Inside” and is assigned to words that are within an entity of a specific type (TYPE). That is, this tag is used when a word is part of a specific entity and indicates the type of that entity. For example, ”I-PER” indicates that the word is part of a person’s name (Person)., and the B-TYPE  [^5] The B in B-TYPE means ”Beginning” and is the tag given to the first word of the second entity when directly followed by an entity of the same type. This tag is used to indicate that a new entity is beginning and to distinguish it from the previous entity. For example, ”B-PER” means that the new person’s name begins with this word. tag is introduced to indicate the start of a new phrase.
B.2 WikiGold:
The Wikigold dataset Balasuriya et al. (2009) is a manually annotated corpus for named entity recognition consisting of a small sample of Wikipedia articles. Words are labeled for each of the four named entity classes (LOC, MISC, ORG, and PER) of CONLL-03 described above.
Appendix C Statistics of Overlapped Entities
LLM-based annotations can overlap different classes on the same entity. We investigate the statistics of overlapped entities.
The fact that the entities overlapping between classes constitute only 521 cases (2.65%) of the CoNLL03 dataset and 109 cases (6.80%) of the WikiGold dataset, and yet the application of label mixing to these specific data points results in a performance improvement across the dataset, is a suggestive observation. This observation implies that the strategic application of label mixing to even a small subset of data can improve the performance balance between entity types.